Rhine river is 160km shorter than St lawerence. The amazon is 800km longer than 5 times the length of the Rhine. Zambezi is is 2x as long as the St lawerence. How long is the Amazon if it is 1440km longer than the sum of the lenghts of the other 3 rivers
Answer by Maths68(1474)

St. Lawerence River = x
Zambezi River = 2x
Rhine River = x-160
Amazon = 800+5(x-160)
=================================================================
Amazon = (St. Lawerence River + Zambezi River + Rhine River)+1440
800+5(x-160) = x+(2x)+(x-160)+1440
800+5x-800= x+2x+x-160+1440
5x=4x+1280
5x= 4x+1280
5x-4x=1280
x=1280
===================================================================
Amazon = 800+5(x-160)
Put the value of x in above equation
Amazon = 800+5(1280-160)
Amazon=800+5(1120)
Amazon = 800+5600
Amazon=6400 km